OpenStudy (zubhanwc3):

for 1/6 dx/dt = x(sqrt(3t^3-1), x(t) give me a moment to write the equation and answers...

OpenStudy (zubhanwc3):

equation is \[\frac{ 1 }{ 6t }\frac{ dx }{ dt }=x \sqrt{3t ^{2}-1}, x(t) =\] answers a) \[e ^{\frac{ 2 }{ 3 }\sqrt{(3t ^{2}-1)^{3}}}\] b) What a was, plus Constant c) \[\frac{ 2 }{ 3 }(3t ^{2}-1)^{\frac{ 2 }{ 3 }}+C\] d) constant * a's answer e) \[\frac{ 2 }{ 3 }\sqrt{3t ^{2}-1}^{3}+C\]

OpenStudy (zubhanwc3):

@ganeshie8 @sourwing

OpenStudy (zubhanwc3):

@zzr0ck3r

OpenStudy (zzr0ck3r):

you need to find \(\int\frac{1}{x}dx=\int6t\sqrt{3t-1}\ dt\)

OpenStudy (zzr0ck3r):

or you can plug in the answers and see which one works...but just solve the integral, its good for you.